# Clifford representatives via the uniform algebraic rank

### v1

@@ Abstract.
MB was supported by CNPq, grant 170255/2023-9. KC was partially supported by NSF FRG grant DMS–2053261. MM is supported by MIUR via the projects PRIN2017SSNZAW (Advances in Moduli Theory and Birational Classification), PRIN 2022L34E7W (Moduli spaces and birational geometry) and PRIN 2020KKWT53 (Curves, Ricci flat Varieties and their Interactions), and is a member of the Centre for Mathematics of the University of Coimbra – UIDB/00324/2020, funded by the Portuguese Government through FCT/MCTES and of the INDAM gorup GNSAGA
@@ 1. Introduction
This inequality can be strict, and much effort has been devoted to describing the gap; see, for example, [Cap13], [AB15], [FJP20], and [AG22] for refinements of the rank and Figure 2 below for a discussion of their relation.
@@
Our main application of these constructions then is the following: A divisor $\underline{d}\in\delta$ is called a Clifford representative if every line bundle $L$ of combinatorial type $\underline{d}$ on any curve $X$ with dual graph $G$ satisfies the Clifford inequality, $r(X,L)\leq\frac{d}{2}$. The following result answers [CLM15, Question 4.7] affirmatively:
@@
Let $\delta$ be an effective divisor class on a graph $G$. Then both the algebraic rank $r^{\operatorname{alg}}(G,\delta)$ and the uniform algebraic rank $r^{\operatorname{ALG}}(G,\delta)$ can be realized by an effective $\underline{d}\in\delta$.
@@ 2.1. Graphs and divisors
Let $G$ be a connected graph of genus $g\geq 2$. We say that $G$ is a stable (respectively semistable) graph if every vertex of weight zero has valency at least $3$ (respectively $2$).
@@
The equivalence class of a divisor $\underline{d}$ is denoted by $[\underline{d}]$. We also use the notation $\delta$ for an element of ${\operatorname{{Pic}}}(G)$, we write $\underline{d}\in\delta$ if $\underline{d}$ is a representative. Since the principal divisors have degree zero, equivalent divisors have the same degree. We set, for an integer $d$,
@@ 2.3. The Baker-Norine combinatorial rank
Let $G$ be a graph possibly with weights or loops, and consider the weightless and loopless graph $G^{\bullet}$ defined by attaching at each vertex $v\in V(G)$ exactly $\omega(v)$ loops based at $v$, and then adding, at each loop, one new vertex in the middle of the edge. For each $\underline{d}\in{\operatorname{{Div}}}(G)$, we can naturally define $\underline{d}^{\bullet}\in{\operatorname{{Div}}}(G^{\bullet})$ by setting $\underline{d}^{\bullet}$ to be 00 at the new vertices of $G^{\bullet}$. In this way we get a natural injective homomorphism $\iota:{\operatorname{{Div}}}(G)\to{\operatorname{{Div}}}(G^{\bullet})$ inducing an injective homomorphism ${\operatorname{{Pic}}}(G)\hookrightarrow{\operatorname{{Pic}}}(G^{\bullet})$. The definition of the Baker-Norine rank has been extended in [AC13] to arbitrary graphs $G$ by setting for any divisor $\underline{d}\in{\operatorname{{Div}}}(G)$:
@@
Notice that the Baker-Norine rank is constant in an equivalence class by definition. We therefore write $r_{G}(\delta)=r_{G}(\underline{d})$, if $\underline{d}$ is a representative of the divisor class $\delta=[\underline{d}]$.
@@ 2.4. The algebraic rank
the variety of isomorphism classes of line bundles of multidegree $\underline{d}$. The ${\operatorname{{Pic}}}^{\underline{d}}(X)$ are the connected components of the degree $d$ Picard variety ${\operatorname{{Pic}}}^{d}(X)$. For every curve $X\in M_{G}$ we have $\underline{\deg}(K_{X})=\underline{k}_{G_{X}}$, where $K_{X}$ denotes the dualizing sheaf on $X$.
In [Cap13] Caporaso introduced a way to give an algebraic interpretation for the Baker-Norine rank of $\underline{\deg}(L)$ on the nodal curves $X$ with dual graph $G$ themselves, i.e., without choosing a smoothing of $X$ as in [Bak08]. More precisely, she defined the so-called algebraic rank $r^{\operatorname{alg}}$, as follows: First, we set
@@ 3.1. Comparison to the algebraic rank
$\underline{d}=(0,3,2)$ $r^{\max}(X,\underline{d})=1$ $r^{\max}(X^{\prime},\underline{d})=2$
$\underline{d}^{\prime}=(3,2,0)$ $r^{\max}(X,\underline{d}^{\prime})=2$ $r^{\max}(X^{\prime},\underline{d}^{\prime})=1$
@@ 3.2. Comparison to the Baker-Norine rank
The inequalities $r^{\operatorname{alg}}(G,\delta)\leq r^{\operatorname{ALG}}(G,\delta)\leq r_{G}(\delta)$ ensure that whenever $r^{\operatorname{alg}}(G,\delta)=r_{G}(\delta)$ also $r^{\operatorname{alg}}(G,\delta)=r^{\operatorname{ALG}}(G,\delta)=r_{G}(\delta)$. This happens, in particular, in the cases summarized in Fact 2.10 (5).
@@ 3.3. Specialization of ranks
More precisely, let $\mathcal{X}\to\operatorname{Spec}R$ be a regular one-parameter smoothing of a nodal curve $X$; that is, a flat family of curves over a discrete valuation ring $R$ with smooth generic fiber $\mathcal{X}_{\eta}$, special fiber $X_{0}$ isomorphic to $X$ and smooth total space. Then any line bundle $\mathcal{L}_{\eta}$ on $\mathcal{X}_{\eta}$ extends to a line bundle $\mathcal{L}$ with central fiber a line bundle $L$ on $X$ whose multidegree we denote by $\underline{d}$. The extension $\mathcal{L}$ is not unique, since twisting by components of $X$ give non-isomorphic extensions of $\mathcal{L}_{\eta}$; but the multidegree of any two extensions lies in the same class $\delta=[\underline{d}]$. In this situation we have:
@@ 3.4. Other notions of rank
where all inequalities can be strict. The first by [Len17, Corollary 3.3], the second by Example 3.4 and for the third cf. Remark 3.9.
@@
In the figure, all entries denote some notion of rank whose value lies between $r(\mathcal{X}_{\eta},\mathcal{L}_{\eta})$ and $r_{G}(\delta)$. Apart from the ones already discussed, these are the rank of metrized complexes $r^{\mathrm{MC}}(\mathcal{C})$ of Amini and Baker [AB15], the refinement of the Baker-Norine rank $r^{\mathrm{FJP}}(\delta)$ given by imposing tropical independence due to Farkas, Jensen and Payne [FJP20, Definition 6.5], and the rank $r^{\mathrm{AG}}(\delta)$ of tropical limit linear series developed by Amini and Gierczak [AG22, §1.5]. [^2] Part of the definition of both $r^{\mathrm{AG}}(\delta)$ and $r^{\mathrm{FJP}}(\delta)$ is the condition imposed for the Baker-Norine rank, in addition to further assumptions. Thus a complete tropical linear series of rank $r$, in their sense, has Baker-Norine rank at most $r$ and $r^{\mathrm{AG}}(\delta),r^{\mathrm{FJP}}(\delta)\leq r_{G}(\delta)$. Two ranks are connected by an arrow if it is known that one is less or equal than the other, with the one towards which the arrow points being bigger. A crossed out line indicates that it is known that no inequality holds in either direction (the result that $r^{\operatorname{alg}}(G,\delta)$ and $r^{\mathrm{MC}}(\mathcal{C})$ are not comparable is due to [Len17, §3], and the examples presented there generalize to $r^{\operatorname{ALG}}(G,\delta)$). Finally, if there is no arrow between two entries, it means that, as far as we are aware, no relation is known.
@@ 4.2. Realization by effective representatives
To prove this statement, we need to recall the Dhar decomposition of the graph $G$ with respect to a subset of vertices $V$. This is a generalization of the Dhar decomposition with respect to a single vertex $v$ (see, e.g., [CLM15, §3.4] for a formulation in our context) studied independently by the first and second author [Bar22, Chr23b].
@@
The kernel of the linear map $H^{0}(X,L)\to H^{0}(Y,L|_{Y})$ is given by global sections $s$ of $L$ that vanish on $Y$. Let $v\not\in V$ be a vertex on which $\underline{d}$ turns negative after a chip-firing move along the complement of $V$. This means that $\underline{d}(v)<v\cdot V=:k$. Since $s$ vanishes by assumption on all components $X_{w}$ corresponding to vertices $w\in V$, $s$ also vanishes on the $k$ points of intersection of $X_{v}$ with the components $X_{w}$. Since $L$ has degree less than $k$ on $X_{v}$, this implies that $s$ vanishes on all of $X_{v}$. Repeating this argument shows that $s$ vanishes on all components $X_{v}$ where $v$ is not in the Dhar set $W_{\mathrm{Dhar}}(\underline{d},V)$. However $W_{\mathrm{Dhar}}(\underline{d},V)=\emptyset$ if $\underline{d}$ is $V$-reduced, since the vertices $v\in W_{\mathrm{Dhar}}(\underline{d},V)$ satisfy by definition $v\cdot W_{\mathrm{Dhar}}(\underline{d},V)^{c}\geq 0$. Thus $s$ needs to vanish on all of $X$ and the kernel of the map $H^{0}(X,L)\to H^{0}(Y,L|_{Y})$ consists only of the zero section. ∎
The point of the next theorem is that in order to calculate the algebraic and the uniform algebraic rank it suffices to restrict to effective divisors $\underline{d}\in\delta$.
@@ 5.1. Existence of Clifford representatives
In [CLM15] the inequality $r^{\operatorname{alg}}(G,\delta)\leq r_{G}(\delta)$ and the fact that $r_{G}$ satisfies the Clifford inequality was used to show that on any nodal curve $X$ with dual graph $G$ there exists a divisor $\underline{d}$ on $G$ such that every line bundle on $X$ of multidegree $\underline{d}$ satisfies the Clifford inequality. Here we are interested in the following stronger notion:
@@
If $G$ is a graph with only one vertex $v$, then $X$ is an irreducible nodal curve and the fact that all line bundles $L$ of degree $d$ on $X$ satisfy the Clifford inequality is shown in the appendix in [EKS88] (actually in loc. cit. the authors show that the Clifford theorem holds more generally for torsion free rank $1$ sheaves on integral curves). The existence of a Clifford representative is also known for weightless and trivalent graphs $G$ by [CLM15, Proposition 4.6].
@@
Let $\delta$ be a divisor class on $G$ of degree $20\leq d\leq 2g-$ 2. Then $\delta$ contains a Clifford representative.
@@ 5.2. Explicit Clifford representatives
In this section, we give an explicit description of Clifford representatives for a large class of graphs. The construction of such Clifford representatives will distinguish between two cases, depending on whether a divisor class is special or not. We begin by introducing the necessary definitions. Recall that we set $\underline{d}^{*}=\underline{k}_{G}-\underline{d}$ with class $\delta^{*}$.
@@
    A divisor $\underline{d}\in{\operatorname{{Div}}}(G)$ is uniform, if both $\underline{d}$ and $\underline{d}^{*}$ are effective.
@@
    A divisor class $\delta\in{\operatorname{{Pic}}}(G)$ is special, if both $\delta$ and $\delta^{*}$ are effective.
@@
$0\leq\underline{d}_{v}\leq 2g_{v}-2+\operatorname{val}(v).$
We have $0>2g_{v}-2+\operatorname{val}(v)$ if and only if $g_{v}=0$ and $\operatorname{val}(v)=1$ (recall that we assume $G$ to be connected). Thus there exist uniform multidegrees on a graph $G$ if and only if $G$ does not contain any vertex $v$ with $g_{v}=0$ and $\operatorname{val}(v)=1$, i.e., $G$ is semistable.
@@
Proposition 5.6 allows us to describe Clifford representatives in special divisor classes, provided that $G$ satisfies the assumptions of the proposition and the class contains a uniform representative. As we mentioned, in general not every special divisor class contains a uniform representative. Our final ingredient is a result that gives a sufficient criterion for the existence of uniform representatives in special classes [Chr23b, Theorem A]. We again state a weaker version adapted to our purposes:
@@
Suppose first $\delta$ is special. Since we assume that for all vertices $v\in V(G)$, if $\omega(v)=0$ then $v$ is adjacent to a loop, $G$ is in particular semistable. Furthermore, by Proposition 5.7 $\delta$ needs to contain a uniform representative $\underline{d}$. By assumption, $G$ is in addition a chain of $2$-edge connected components, and hence any uniform divisor $\underline{d}$ is a Clifford representative by Proposition 5.6.

### v2

@@ Abstract.
MB was supported by CNPq, grant 170255/2023-9. KC was partially supported by NSF FRG grant DMS–2053261 and is member of the INdAM group GNSAGA. MM is supported by MIUR via the projects PRIN2017SSNZAW (Advances in Moduli Theory and Birational Classification), PRIN 2022L34E7W (Moduli spaces and birational geometry) and PRIN 2020KKWT53 (Curves, Ricci flat Varieties and their Interactions), and is a member of the Centre for Mathematics of the University of Coimbra – UIDB/00324/2020, funded by the Portuguese Government through FCT/MCTES and of the INdAM group GNSAGA
2020 Mathematics Subject Classification: Primary – 14H51; Secondary – 14T20
@@ 1. Introduction
This inequality can be strict, and much effort has been devoted to describing the gap; see, for example, [Cap13], [AB15], [FJP20], and [AG22] for refinements of the rank and Figure 2 in Section 3.4 below for a discussion of their relation.
@@
Our main application of these constructions is the following: A divisor $\underline{d}\in\delta$ is called a Clifford representative if every line bundle $L$ of combinatorial type $\underline{d}$ on any curve $X$ with dual graph $G$ satisfies the Clifford inequality, $r(X,L)\leq\frac{d}{2}$. The following result answers [CLM15, Question 4.7] affirmatively:
@@
Let $\delta$ be an effective divisor class on a graph $G$. Then both the algebraic rank $r^{\operatorname{alg}}(G,\delta)$ and the uniform algebraic rank $r^{\operatorname{ALG}}(G,\delta)$ can be realized by an effective representative $\underline{d}\in\delta$.
@@ 2.1. Graphs and divisors
Let $G$ be a connected graph of genus $g\geq 2$. We say that $G$ is a stable (respectively semistable) graph if every vertex of weight zero has valence at least $3$ (respectively $2$).
@@
The equivalence class of a divisor $\underline{d}$ is denoted by $[\underline{d}]$. We also use the notation $\delta$ for an element of ${\operatorname{{Pic}}}(G)$, and we write $\underline{d}\in\delta$ if $\underline{d}$ is a representative. Since the principal divisors have degree zero, equivalent divisors have the same degree. We set, for an integer $d$,
@@ 2.3. The Baker-Norine combinatorial rank
Let $G$ be a graph possibly with weights or loops, and consider the weightless and loopless graph $G^{\bullet}$ defined by attaching at each vertex $v\in V(G)$ exactly $\omega(v)$ loops based at $v$, and then adding, at each loop, one new vertex subdividing the edge. For each $\underline{d}\in{\operatorname{{Div}}}(G)$, we can naturally define $\underline{d}^{\bullet}\in{\operatorname{{Div}}}(G^{\bullet})$ by setting $\underline{d}^{\bullet}$ to be 00 at the new vertices of $G^{\bullet}$. In this way we get a natural injective homomorphism $\iota:{\operatorname{{Div}}}(G)\to{\operatorname{{Div}}}(G^{\bullet})$ inducing an injective homomorphism ${\operatorname{{Pic}}}(G)\hookrightarrow{\operatorname{{Pic}}}(G^{\bullet})$. The definition of the Baker-Norine rank has been extended in [AC13] to arbitrary graphs $G$ by setting for any divisor $\underline{d}\in{\operatorname{{Div}}}(G)$:
@@
Notice that the Baker-Norine rank is constant in an equivalence class by definition. We therefore write $r_{G}(\delta)=r_{G}(\underline{d})$ if $\underline{d}$ is a representative of the divisor class $\delta=[\underline{d}]$.
@@ 2.4. The algebraic rank
the variety of isomorphism classes of line bundles of multidegree $\underline{d}$. The varieties ${\operatorname{{Pic}}}^{\underline{d}}(X)$ are the connected components of the degree $d$ Picard variety ${\operatorname{{Pic}}}^{d}(X)$. For every curve $X\in M_{G}$, we have $\underline{\deg}(K_{X})=\underline{k}_{G_{X}}$, where $K_{X}$ denotes the dualizing bundle on $X$.
In [Cap13], Caporaso introduced a way to give an algebraic interpretation for the Baker-Norine rank of $\underline{\deg}(L)$ on the nodal curves $X$ with dual graph $G$ themselves, i.e., without choosing a smoothing of $X$ as in [Bak08]. More precisely, she defined the so-called algebraic rank $r^{\operatorname{alg}}$, as follows: First, we set
@@ 3.1. Comparison to the algebraic rank
$\underline{d}=(0,3,2)$ $r^{\max}(X,\underline{d})=1$ $r^{\max}(X^{*},\underline{d})=2$
$\underline{d}^{\prime}=(3,2,0)$ $r^{\max}(X,\underline{d}^{\prime})=2$ $r^{\max}(X^{*},\underline{d}^{\prime})=1$
@@ 3.2. Comparison to the Baker-Norine rank
The inequalities $r^{\operatorname{alg}}(G,\delta)\leq r^{\operatorname{ALG}}(G,\delta)\leq r_{G}(\delta)$ ensure that whenever $r^{\operatorname{alg}}(G,\delta)=r_{G}(\delta)$, also $r^{\operatorname{alg}}(G,\delta)=r^{\operatorname{ALG}}(G,\delta)=r_{G}(\delta)$. This happens, in particular, in the cases summarized in Fact 2.10 (5).
@@ 3.3. Specialization of ranks
More precisely, let $\mathcal{X}\to\operatorname{Spec}R$ be a regular one-parameter smoothing of a nodal curve $X$; that is, a flat family of curves over a discrete valuation ring $R$ with smooth generic fiber $\mathcal{X}_{\eta}$, special fiber $X_{0}$ isomorphic to $X$ and smooth total space. Then any line bundle $\mathcal{L}_{\eta}$ on $\mathcal{X}_{\eta}$ extends to a line bundle $\mathcal{L}$ with central fiber a line bundle $L$ on $X$ whose multidegree we denote by $\underline{d}$. The extension $\mathcal{L}$ is not unique, since twisting by components of $X$ gives non-isomorphic extensions of $\mathcal{L}_{\eta}$; but the multidegree of any two extensions lies in the same class $\delta=[\underline{d}]$. In this situation we have:
@@ 3.4. Other notions of rank
where all inequalities can be strict. The first by [Len17, Corollary 3.3], the second by Example 3.4 and for the third cf. Remark 3.9. Note, however, that in all examples mentioned only one of the above inequalities is strict. We do not know whether it is possible to find examples in which (at least) two inequalities are strict at the same time.
@@
In the figure, all entries denote some notion of rank whose value lies between $r(\mathcal{X}_{\eta},\mathcal{L}_{\eta})$ and $r_{G}(\delta)$. Apart from the ones already discussed, these are the ranks of metrized complexes $r^{\mathrm{MC}}(\mathcal{C})$ of Amini and Baker [AB15], the refinement of the Baker-Norine rank $r^{\mathrm{FJP}}(\delta)$ given by imposing tropical independence due to Farkas, Jensen and Payne [FJP20, Definition 6.5], and the rank $r^{\mathrm{AG}}(\delta)$ of tropical limit linear series developed by Amini and Gierczak [AG22, §1.5]. [^2] Part of the definition of both $r^{\mathrm{AG}}(\delta)$ and $r^{\mathrm{FJP}}(\delta)$ is the condition imposed for the Baker-Norine rank, in addition to further assumptions. Thus a complete tropical linear series of rank $r$, in their sense, has Baker-Norine rank at least $r$ and $r^{\mathrm{AG}}(\delta),r^{\mathrm{FJP}}(\delta)\leq r_{G}(\delta)$. Two ranks are connected by an arrow if it is known that one is less or equal than the other, with the one towards which the arrow points being bigger. A crossed-out line indicates that it is known that no inequality holds in either direction (the result that $r^{\operatorname{alg}}(G,\delta)$ and $r^{\mathrm{MC}}(\mathcal{C})$ are not comparable is due to [Len17, §3], and the examples presented there generalize to $r^{\operatorname{ALG}}(G,\delta)$). Finally, if there is no arrow between two entries, it means that, as far as we are aware, no relation is known.
@@ 4.2. Realization by effective representatives
To prove this statement, we need to recall the Dhar decomposition of the graph $G$ with respect to a subset of vertices $V$. This is a generalization of the Dhar decomposition with respect to a single vertex $v$ (see, e.g., [CLM15, §3.4] for a formulation in our context) studied independently by the first and second authors [Bar22, Chr23b].
@@
The kernel of the linear map $H^{0}(X,L)\to H^{0}(Y,L|_{Y})$ is given by global sections $s$ of $L$ that vanish on $Y$. Let $v\not\in V$ be a vertex on which $\underline{d}$ turns negative after a chip-firing move along the complement of $V$. This means that $\underline{d}(v)<v\cdot V=:k$. Since $s$ vanishes by assumption on all components $X_{w}$ corresponding to vertices $w\in V$, $s$ also vanishes on the $k$ points of intersection of $X_{v}$ with the components $X_{w}$. Since $L$ has degree less than $k$ on $X_{v}$, this implies that $s$ vanishes on all of $X_{v}$. Repeating this argument shows that $s$ vanishes on all components $X_{v}$ where $v$ is not in the Dhar set $W_{\mathrm{Dhar}}(\underline{d},V)$. However, $W_{\mathrm{Dhar}}(\underline{d},V)=\emptyset$ if $\underline{d}$ is $V$-reduced, since the vertices $v\in W_{\mathrm{Dhar}}(\underline{d},V)$ satisfy by definition $v\cdot W_{\mathrm{Dhar}}(\underline{d},V)^{c}\geq 0$. Thus $s$ needs to vanish on all of $X$, and the kernel of the map $H^{0}(X,L)\to H^{0}(Y,L|_{Y})$ consists only of the zero section. ∎
The point of the next theorem is that in order to calculate the algebraic and the uniform algebraic rank, it suffices to restrict to effective divisors $\underline{d}\in\delta$.
@@ 5.1. Existence of Clifford representatives
In [CLM15] the inequality $r^{\operatorname{alg}}(G,\delta)\leq r_{G}(\delta)$ and the fact that $r_{G}$ satisfies the Clifford inequality were used to show that on any nodal curve $X$ with dual graph $G$ there exists a divisor $\underline{d}$ on $G$ such that every line bundle on $X$ of multidegree $\underline{d}$ satisfies the Clifford inequality. Here we are interested in the following stronger notion:
@@
If $G$ is a graph with only one vertex $v$, then $X$ is an irreducible nodal curve and the fact that all line bundles $L$ of degree $d$ on $X$ satisfy the Clifford inequality is shown in the appendix in [EKS88] (actually in loc. cit. the authors show that the Clifford theorem holds more generally for torsion-free rank $1$ sheaves on integral curves). The existence of a Clifford representative is also known for weightless and trivalent graphs $G$ by [CLM15, Proposition 4.6].
@@
Let $\delta$ be a divisor class of degree $20\leq d\leq 2g-$ 2 on a graph $G$ of genus $g$. Then $\delta$ contains a Clifford representative.
@@ 5.2. Explicit Clifford representatives
In this section, we give an explicit description of Clifford representatives for a large class of graphs. The construction of such Clifford representatives will distinguish between two cases, depending on whether a divisor class is special or not. We begin by introducing the necessary definitions. Recall that we set $\underline{d}^{*}=\underline{k}_{G}-\underline{d}$ and that we denote by $\delta^{*}$ the class of $\underline{d}^{*}$.
@@
    A divisor $\underline{d}\in{\operatorname{{Div}}}(G)$ is uniform if both $\underline{d}$ and $\underline{d}^{*}$ are effective.
@@
    A divisor class $\delta\in{\operatorname{{Pic}}}(G)$ is special if both $\delta$ and $\delta^{*}$ are effective.
@@
$0\leq\underline{d}_{v}\leq 2\omega(v)-2+\operatorname{val}(v).$
We have $0>2\omega(v)-2+\operatorname{val}(v)$ if and only if $\omega(v)=0$ and $\operatorname{val}(v)=1$ (recall that we assume $G$ to be connected). Thus there exist uniform multidegrees on a graph $G$ if and only if $G$ does not contain any vertex $v$ with $\omega(v)=0$ and $\operatorname{val}(v)=1$, i.e., $G$ is semistable.
@@
Proposition 5.6 allows us to describe Clifford representatives in special divisor classes, provided that $G$ satisfies the assumptions of the proposition and the class contains a uniform representative. In general, not every special divisor class contains a uniform representative. Our final ingredient is a result that gives a sufficient criterion for the existence of uniform representatives in special divisor classes [Chr23b, Theorem B]. We again state a weaker version adapted to our purposes:
@@
Suppose first that $\delta$ is special. Since we assume that for all vertices $v\in V(G)$, if $\omega(v)=0$ then $v$ is adjacent to a loop, $G$ is in particular semistable. Furthermore, by Proposition 5.7, $\delta$ needs to contain a uniform representative $\underline{d}$. By assumption, $G$ is in addition a chain of $2$-edge connected components, and hence any uniform divisor $\underline{d}$ is a Clifford representative by Proposition 5.6.